# Quantifying Emotional Arousal through Pupillary Response: A Novel Approach for Isolating the Luminosity Effect and Predicting Affective States

### v1

@@ II-B2 Arousal Ground Truth
As ground truth for the ADM, we used the participants’ questionnaire responses, mapping their reported emotion scales into the arousal/valence two-dimensional space. Initially, we explored factor analysis for this purpose; however, this failed to distribute emotions effectively across valence and arousal dimensions, resulting in suboptimal model training outcomes (see results VII). Instead, we employed Individual Scaling (INDSCAL) [kim2016identifying], a multidimensional scaling approach that reduces high-dimensional data into interpretable lower-dimensional representations. This approach condensed participants’ responses from 12 emotional dimensions into two key dimensions: valence (x-coordinate) and arousal (y-coordinate) (see Figure 5). INDSCAL creates a common multidimensional space, like multidimensional scaling, and an individual space for each subject. This methodology improved the precision of predicting emotional states by providing a robust ground truth for training the ADM.
II-B3 Pupil Data Pre-processing
@@
II-B4 Arousal Effect Prediction
@@
We trained the model described by the Equations 5 and 6 by calculating , , , and frame by frame and the coefficients , , , , , and for each video clip. We hypothesized that since there was also an arousal component (in addition to the light component), the fitting would be worse and that the error in the fitting would be due specifically to the effect of arousal, given that the model was designed to capture only the component due to luminosity. In other words, we hypothesized that the measured pupil size would be equal to the pupil size predicted by the LEPM model as an effect of luminosity plus a residual:
@@ II-B5 ADM Testing
When completing the questionnaire at the end of each video clip, participants were asked to report their emotion and arousal levels without specifying the exact moment in the video when they occurred. To identify the emotionally salient intervals that would allow us to predict arousal accurately at each relevant moment throughout the video duration, we randomly selected a sub-sample of 10 participants from the 55 who previously participated in the study. Participants were asked to indicate the salient intervals of each video. The final salient interval of a video was defined as the interval for which there was the highest degree of agreement among the 10 participants. The overlap between the chosen interval and that indicated by each participant was, on average, 74% 15%, considering all the videos, indicating a reasonable degree of agreement. For some videos, e.g., boring ones, the salient intervals corresponded to the entire video, while for other videos, e.g., scary ones, only some segments were salient. Finally, we only computed the average pupil size for each video and participant across the salient intervals, with and without our luminosity correction.
@@
We then calculated the pupil size corrected for luminosity and self-reported arousal for each video and each participant remaining in our dataset. We then assumed a simple linear relationship between the size of the pupil corrected for luminosity (component due to arousal) and self-reported arousal:
We fitted the model 12 to all the videos and all the participants (except one). We obtained a good fit (see Results).
@@
We then repeated the same procedure for pupil size not corrected for luminosity. In this case, in Equations 12 and 13, must be replaced with . We predicted the self-reported arousal value for the eliminated participant starting from the pupil size, using Equation 13, and for all the videos. We compared the obtained values with the ground truth for the luminosity-corrected and uncorrected pupil size. We repeated this procedure, eliminating one participant at a time, and finally calculated the average results for our sample.
@@
All the code relative to the models presented in this paper was written in Python.
III Results
III-A Results of the Luminosity Effect Prediction Model (LEPM)
@@
The model illustrated in section “A. Development of Luminosity Effect Prediction Model” effectively predicts pupil size based on the RGB intensity values of the images on the screen in both dark and bright environments. The calibration procedure described above ensures that the method is flexible and adapted to inter-subjective differences, as well as the settings and type of screen used. As explained in section II-A, we initially used two different approaches, named “gray-based” and the “color-based”, and then we used a method consisting of combining both approaches, named the “combined” approach. We assessed each approach using leave-one-image-out cross-validation, i.e., training the model on 27 images, eliminating one image at a time, and predicting the pupil size measured when the eliminated image was shown.
@@
The results obtained from the 18 participants in a dark laboratory, i.e., the mean Pearson correlation coefficient with the mean and max p-value, the mean R2, the mean NRMSE777Normalized Root Mean Squared Error (NRMSE) is a scale-independent error metric that allows comparison between datasets or models with different scales. We computed it by dividing the Root Mean Squared Error (RMSE) by a normalization factor, consisting of the range (max-min) of the observed data., and the mean percentage error is shown in Table II.
@@
Table III, on the other hand, shows the results obtained by aggregating the data across 18 participants.
| Method | Correlation | R2 Score | NRMSE | Average Error |
@@
| Method | Correlation | R2 Score | NRMSE | Average Error |
@@
The results for the “combined” method were better than when using only the “gray-based” method or only the “color-based” method, as shown in tables II and III, demonstrating that the two methods contribute synergistically to the evaluation of pupil size. Since the best results were achieved with the “combined” method, all the subsequent results were obtained using this approach.
As explained in the paragraph Testing the Luminosity Effect Prediction Model (LEPM) in a well-lit laboratory and with non-monochrome images, we tested the LEPM and calibration procedure on non-monochrome, non-primary color images and in both a dark and a lit environment. The test was conducted with 10 participants using a video composed of 27 monochrome images and 46 non-monochrome emotionally neutral images with different colors and brightness.
Figures 6 and 8 show the results for participant IFL3 in a dark and a well-lit laboratory, respectively. The average results across participants, obtained in the dark and well-lit laboratories, are shown in Table IV.
Figures 7 and 9 and Table V, instead, show the results obtained by aggregating the data from all the participants.
| Method | Correlation | R2 Score | NRMSE | Average Error |
@@
| Well-light laboratory | 0.76 0.045 (mean p , max p | 0.58 0.680 | 0.14 0.014 | 7.72% 1.66% |
| Method | Correlation | R2 Score | NRMSE | Average Error |
@@
| Well-light laboratory | 0.82 (mean p , max p | 0.68 | 0.11 | 7.50% 0.83% |
The average error between the two conditions is about 0.2%, and the difference in NRMSE is about 0.02, highlighting the model’s effectiveness across dark and well-lit settings. Additionally, the model performs consistently across various monitor brightness and contrast settings and under different ambient lighting levels. Therefore, there was no need to modify the model to account for changes in environmental luminosity, given that our calibration procedure takes this into account effectively. However, we repeated the calibration every 20 minutes for very long experimental sessions.
III-B Results of the Arousal Detection Model (ADM)
To test the ADM, we predicted the arousal of a participant watching video clips based on the measured pupil size and compared it with the ground truth. For each participant and each video clip, we plotted the value of the measured pupil size throughout the video clip. This is shown in Figure 10 (green line) for a particular participant, and a video with increasing high emotional intensity (Figure 10(a), and one with low intensity (Figure 10(b)). In the same figure, we show the average RGB value frame by frame, representing the luminous intensity of the video (red line). We then calculated the pupil size component due to luminosity using the LEPM (see Figure 10, blue line). Finally, we calculated the pupil size component due to arousal by subtracting from , according to Equation 11 (see the black line in Figure 10).
@@ III-B Results of the Arousal Detection Model (ADM)
We then calculated the average of the (black line in Figure 10) in the salient intervals, as explained in the paragraph Model Testing of section II-B. For example, for the videos represented in Figure 10, the salient intervals were [0s, 20s] for video clip (a) and [0s, 10s] for video clip (b). Finally, we plotted the pupil size corrected for luminosity versus the self-reported arousal for all video clips and each participant, which is shown in Figure 11 for participant XPi3pA, where the red circles correspond to each video clip (see figure 11 for more information). For that participant, we obtained a correlation of 0.71 (p = 1.684e-06); see the red line. For the measured pupil size, non-corrected for luminosity , the correlation with the ground truth arousal was much worse, e.g., see for participant XPi3pA in Figure 11, where the blue dots correspond to each video clip. We obtained a correlation of 0.01 (p = 0.971) for that participant (see blue line in the figure 11).
Table VI shows all participants’ average results. Correcting for luminosity dramatically increased correlation compared to non-corrected pupil size. The most surprising result, however, is that pupil size had no predictive power for arousal without correcting for luminosity since its correlation with self-reported arousal was not significantly different from zero (mean p = 0.2283).
@@
To further test the model performance, we computed the predicted arousal utilizing the equations 12 and 13 and using a leave-one-participant out cross-validation as explained in the paragraph ADM Testing. We plotted the predicted arousal against the ground truth one both with and without correction for luminosity, as shown in Figure 12 for the participant XPI3pA.
The average results across all the participants are shown in the left panel in Table VII. Again, correcting for luminosity produces a dramatic increase in correlation compared to uncorrected pupil size. While the left panel shows the results obtained by calculating self-reported arousal using individual scaling, the right panel shows the results obtained by calculating self-reported arousal using factor analysis. As mentioned in the paragraph Development of Arousal Detection Model, INDSCAL works much better than factor analysis in this case.
@@
Finally, we noticed that, in the case of pupil size corrected for luminosity, the average values of the coefficients of Equation 12 were = 0.3463 0.0551, = -0.0126 0.00038 and the fit was quite good ( = 0.567 0.073). The variance of the coefficients and is since leaving one participant out at a time changes the slope and the intercept of the regression line. However, this variation is minuscule (the standard deviation is much smaller than the mean), indicating that our sample was sufficiently large. In the case of pupil size not corrected for luminosity, the average values of the coefficients were = 0.36723 0.06341, = 3.8296 0.0120, and the fit was inferior ( = 0.023 0.007).
III-C Testing models developed by other researchers with our data
Testing Nakayama’s Hyperbolic Model with our data
Nakayama and colleagues [nakayama2021controlling] found a hyperbolic relationship between luminosity and pupil size, which they called standardized pupil size:
where is the luminosity.
The function should predict the variation in pupil size with luminosity to within a constant multiplicative factor. The authors, therefore, hypothesized that if they use emotionally neutral images, dividing the measured pupil size by the function should yield a constant factor that they called the “compensated pupil size”. If, on the other hand, images with emotional content are used, a variation is found concerning the constant value, and the variation is due to the arousal. We have tested this model on our 73 emotionally neutral test images and obtained a coefficient of variation of 0.09 0.025. Our exponential model works better, obtaining a coefficient of variation of 0.03 0.006.
We then used Nakayama’s Hyperbolic Model with our emotional video clips. We used the calibration procedure to calculate the only coefficient of the model, consisting of the ratio between the measured pupil size and the standardized pupil size. The comparison with our method is presented in the table VIII. Nakayama’s method worked less well than ours. For example, there is a worsening of the NRMSE of 51.8%.
Testing Linear Models (e.g., Raiturkar’s and Asano’s Model) with our data
Both Raiturkar et al. [raiturkar2016decoupling] and Asano et al. [asano2021neural, asano2021pupil] have developed dynamic linear models that take into account the temporal response of pupil dilation. We did not develop dynamic models because we took the average pupil size over relatively long intervals (at least 5 seconds). However, we wanted to test linear models with our data and obtained very poor results, as shown in the table VIII.
Testing our model without self-reported arousal
@@ III-C Testing models developed by other researchers with our data
As mentioned, where possible, we applied to our data the methods developed by other researchers to exploit the fact that we had recorded the self-reported arousal for each participant and each video, something that, to the best of our knowledge, no other research on the effect of luminosity on pupil size measurement has done. We investigated how important it is to include self-reported arousal. To do this, we used the approach other researchers have used: asking 10 independent judges (new participants) to give an arousal value to each video (e.g., Raiturkar et al. [raiturkar2016decoupling]). For each clip, we took the average value. We then used that arousal value in our method, instead of the self-reported one, to observe how much the average arousal of the independent judges could be predicted by pupil size corrected for the luminosity of our 47 participants. The comparison between the results obtained with our model using as arousal values those self-reported by the 47 participants, i.e., what we described so far, and the results obtained with our model using as arousal values those given by the 10 independent judges, is shown in table VIII. The model trained on the arousal reported by the 10 independent judges provides clearly worse results than the one trained on the arousal reported by the 47 participants. For example, there is a worsening in NRMSE of 188%. This is because when training a model, one must use a ground truth recorded from the same subjects from which the predictors are recorded, in this case, respectively, the self-reported arousal and pupil size of the 47 participants in the study.
@@
| Metric | Our Model | Hyperbolic Model [nakayama2021controlling] | Linear Model [raiturkar2016decoupling, asano2021neural] | Our Model (without self-reported arousal) [raiturkar2016decoupling] |

### v2

@@ 2.2.2 Arousal Ground Truth
As ground truth for the ADM, we used the participants’ questionnaire responses, mapping their reported emotion scales into the arousal/valence two-dimensional space. Initially, we explored factor analysis for this purpose; however, this failed to distribute emotions effectively across valence and arousal dimensions, resulting in suboptimal model training outcomes (see results 7). Instead, we employed Individual Scaling (INDSCAL) Kim et al. (2016), a multidimensional scaling approach that reduces high-dimensional data into interpretable lower-dimensional representations. This approach condensed participants’ responses from 12 emotional dimensions into two key dimensions: valence (x-coordinate) and arousal (y-coordinate) (see Figure 5). INDSCAL creates a common multidimensional space, like multidimensional scaling, and an individual space for each subject. This methodology improved the precision of predicting emotional states by providing a robust ground truth for training the ADM.
2.2.3 Pupil Data Pre-processing
@@
2.2.4 Arousal Effect Prediction
@@
We trained the model described by the Equations 7 and 8 by calculating , , , and frame by frame and the coefficients , , , , , and for each video clip. We hypothesized that since there was also an arousal component (in addition to the light component), the fitting would be worse and that the error in the fitting would be due specifically to the effect of arousal, given that the model was designed to capture only the component due to luminosity. In other words, we hypothesized that the measured pupil size would be equal to the pupil size predicted by the LEPM model as an effect of luminosity plus a residual:
@@ 2.2.5 ADM Testing
When completing the questionnaire at the end of each video clip, participants were asked to report their emotion and arousal levels without specifying the exact moment in the video when they occurred. To identify the emotionally salient intervals, and would allow us to predict arousal accurately at each relevant moment throughout the video duration, we randomly selected a sub-sample of 10 participants from the 55 who previously participated in the study. Participants were asked to indicate the salient intervals of each video. The final salient interval of a video was defined as the interval for which there was the highest degree of agreement among the 10 participants. The overlap between the chosen interval and that indicated by each participant was, on average, 74% 15%, considering all the videos, indicating a reasonable degree of agreement. For some videos, e.g., boring ones, the salient intervals corresponded to the entire video, while for other videos, e.g., scary ones, only some segments were salient. Finally, we only computed the average pupil size for each video and participant across the salient intervals, with and without our luminosity correction.
@@
We then calculated the pupil size corrected for luminosity and self-reported arousal
for each video and each participant remaining in our dataset. We then assumed a simple linear relationship between the size of the pupil corrected for luminosity (component due to arousal) and self-reported arousal:
We fitted the model 14 to all the videos and all the participants (except one). We obtained a good fit (see Results). As mentioned in the introduction, it has recently been discovered that the interaction between the arousal effect and the effect of light on pupil size is multiplicative. However, we assumed, as a first approximation, that for relatively small variations in luminosity, a linear approximation could be used.
@@
We then repeated the same procedure for pupil size not corrected for luminosity. In this case, in Equations 14 and 15, must be replaced with . We predicted the self-reported arousal value for the eliminated participant starting from the pupil size, using Equation 15, and for all the videos. We compared the obtained values with the ground truth for the
luminosity-corrected and uncorrected pupil size. We repeated this procedure, eliminating one participant at a time, and finally calculated the average results for our sample.
2.3 Development of the Gradient Boost Regression-based Self-reported Arousal Prediction Model (GBR-SAPM)
We developed a relatively complex model for the first phase (LEPM) and a very simple one (ADM) for the second phase. As noted above, the ADM was deliberately designed to be simple in order to demonstrate that if a researcher simply subtracts the component , obtained through the LEPM, from the measured pupil size, they obtain a significant improvement in assessing the experienced emotional arousal– without resorting to further pre-processing or machine learning techniques (see Results).
Nevertheless, when more advanced pre-processing and machine learning methods are applied, predictive performance improves further, as we demonstrate in this section.
We first calculated the signal, using the LEPM model, as described in Equation 13. Then, from each of the three signals , , and , we extracted the following features:
Statistical Moments: Mean, variance, skewness, and kurtosis of the pupil diameter. These metrics provide a summary of the overall activation level and the distribution of the pupil response over the stimulus duration Bradley et al. (2008).
Dynamic Features: The first and second derivatives of the pupil response. These features capture the temporal dynamics of dilation, which are critical for distinguishing the slower, sympathetic-driven arousal response from faster, reflex-driven changes Cacioppo et al. (2007).
The decision to extract features from all three of the above-mentioned signals, rather than using only, stems from the recent discovery that the relationship between total pupil size, the part driven by emotional arousal, and the part driven by luminosity, is not additive but involves non-linear interactionsPan et al. (2022).
We then used a Gradient Boosting regressor (XGBoost), which was preferred over traditional linear models due to its ability to model complex, non-linear interactions. XGBoost is particularly robust against multicollinearity, a common problem in physiological signal processing, and uses regularization (L1/L2) to prevent overfitting.
To ensure the model generalizes to unseen subjects, we implemented a nested Leave-One-Participant-Out (LOPO) cross-validation framework:
A.
Outer Loop (Generalization): The model is trained on participants and tested on the held-out participant. This mimics a real-world scenario where the system encounters a new user without prior calibration Leinonen et al. (2024).
B.
Inner Loop (Optimization): Inside each training set, we performed a five-fold cross-validation to optimize hyper parameters (e.g., learning rate, max depth).
We evaluated the model’s ability to track subjective arousal using the Coefficient of Determination (), Pearson correlation coefficient (), and Normalized Root Mean Square Error ().
We also trained the model to predict valence, although we expected worse results than for arousal. While there is consensus in the literature that pupil size is a reliable indicator of arousal, its relationship with valence is much more controversial. There are some studies, however, such as that by Tarnowski et al. Tarnowski et al. (2020), in which machine learning (SVM and neural networks) was used to classify six basic emotions (happiness, sadness, fear, etc.), which inherently implies the distinction between positive and negative valence. Inspired by this study, we trained our model to also predict emotional valence, and the training was done completely independently from the training done to predict emotional arousal.
@@ 2.3 Development of the Gradient Boost Regression-based Self-reported Arousal Prediction Model (GBR-SAPM)
All the code relative to the models presented in this paper was written in Python with libraries like pandas, numpy, scikit-learn, matplotlib, seaborn, sklearn, xgboost, and os.
3 Results
3.1 Results of the Luminosity Effect Prediction Model (LEPM)
@@
The model illustrated in section “A. Development of Luminosity Effect Prediction Model” effectively predicts pupil size based on the RGB intensity values of the images on the screen in both dark and bright environments. The calibration procedure described above ensures that the method is flexible and adapted to inter-subjective differences, as well as the settings and type of screen used. As explained in section 2.1, we initially used two different approaches, named “gray-based” and the “color-based”, and then we used a method consisting of combining both approaches, named the “combined” approach. We assessed each approach using leave-one-image-out cross-validation, i.e., training the model on 27 images, eliminating one image at a time, and predicting the pupil size measured when the eliminated image was shown.
@@
The results obtained from the 18 participants tested in dark laboratory conditions – mean Pearson correlation coefficient with mean and max p-value, mean , mean 777Normalized Root Mean Squared Error () is a scale-independent error metric that allows comparison between datasets or models with different scales. We computed it by dividing the Root Mean Squared Error (RMSE) by a normalization factor, consisting of the range (max-min) of the observed data. and mean percentage error – are shown in Table 2.
@@
Table 3, on the other hand, shows the results obtained by combining the data from all 18 participants as if they were a single participant.
| Method | Correlation | Score | Average Error | |
@@
| Method | Correlation | Score | Average Error | |
@@
The results for the “combined” method were better than when using only the “gray-based” method or only the “color-based” method, as shown in tables 2 and 3, demonstrating that the two methods contribute synergistically to the evaluation of pupil size. Since the best results were achieved with the “combined” method, all the subsequent results were obtained using this approach.
As explained in the paragraph Testing the Luminosity Effect Prediction Model (LEPM) in
a normal laboratory conditions and with non-monochrome images, we tested the LEPM and calibration procedure on non-monochrome, non-primary color images and in both a dark and a lit environment. The test was conducted with 10 participants using a video composed of 27
monochrome images and 46 non-monochrome emotionally neutral images with different colors and brightness.
Figures 6 and 8 show the results for participant IFL3 in a dark and a normal laboratory conditions, respectively. The average results across participants, obtained in the dark and normal laboratory conditions, are shown in Table 4.
Figures 7 and 9 and Table 5, instead, show the results obtained by aggregating the data from all the participants.
| Method | Correlation | Score | Average Error | |
@@
| normal laboratory conditions | 0.76 0.045 (mean p , max p | 0.58 0.0680 | 0.14 0.014 | 7.72% 1.66% |
| Method | Correlation | Score | Average Error | |
@@
| normal laboratory conditions | 0.82 (mean p , max p | 0.68 | 0.11 | 7.50% 0.83% |
The average error between the two conditions is about 0.2%, and the difference in is about 0.02, highlighting the model’s effectiveness across dark and normal laboratory conditions settings. Additionally, the model performs consistently across various monitor brightness and contrast settings and under different ambient lighting levels. Therefore, there was no need to modify the model to account for changes in environmental luminosity, given that our calibration procedure takes this into account effectively. However, we repeated the calibration every 20 minutes for very long experimental sessions.
3.2 Results of the Arousal Detection Model (ADM)
To test the ADM, we predicted the arousal of participants watching video clips and compared it with the ground truth (actual pupil size). For each participant and each video clip, we plotted the value of the measured pupil size throughout the video clip. This is shown in Figure 10 (green line) for a particular participant, and a video with increasing high emotional intensity (Figure 10(a), and one with low intensity (Figure 10(b)). In the same figure, we show the average RGB value frame by frame, representing the luminous intensity of the video (red line). We then calculated the pupil size component due to luminosity using the LEPM (see Figure 10, blue line). Finally, we calculated the pupil size component due to arousal by subtracting from , according to Equation 13 (see the black line in Figure 10).
@@ 3.2 Results of the Arousal Detection Model (ADM)
We then calculated the average of the (black line in Figure 10) in the salient intervals, as explained in the paragraph Model Testing of section 2.2. For example, for the videos represented in Figure 10, the salient intervals were [0s, 20s] for video clip (a) and [0s, 10s] for video clip (b). Finally, we plotted the pupil size corrected for luminosity versus the self-reported arousal for all video clips and each participant, which is shown in Figure 11 for participant XPi3pA, where the red circles correspond to each video clip (see figure 11 for more information). For that participant, we obtained a correlation of 0.71 (; see the red line). For the measured pupil size, non-corrected for luminosity , the correlation with the ground truth arousal was much worse – see participant XPi3pA in Figure 11, where the blue dots correspond to each video. clip. We obtained a correlation of 0.01 (p = 0.971) for that participant (see blue line in the figure 11).
Table 6 shows the average of all participants. Correcting pupil size for luminosity dramatically increased correlation compared to non-correcting it. The most surprising result, however, was that pupil size had no predictive power for arousal without correcting for luminosity, as shown by the correlation with self-reported arousal not being significantly different from zero (mean p = 0.2283).
@@
To further test the model performance, we computed the predicted utilizing the equations 14 and 15 and using a leave-one-participant out cross-validation as explained in the paragraph ADM Testing. We plotted the predicted arousal against the ground truth one both with and without correction for luminosity, as shown in Figure 12 for participant XPI3pA.
The average results across all participants are shown in the left panel in Table 7. Again, correcting for luminosity produces a dramatic increase in correlation compared to uncorrected pupil size. While the left panel shows the results obtained by calculating self-reported arousal using individual scaling, the right panel shows the results obtained by calculating self-reported arousal using factor analysis. As mentioned in the paragraph Development of Arousal Detection Model, INDSCAL works much better than factor analysis in this case.
@@
Finally, we noticed that, in the case of pupil size corrected for luminosity, the average values of the coefficients of Equation 14 were = 0.3463 0.0551, = -0.0126 0.00038 and the fit was quite good ( = 0.567 0.073). The variance of the coefficients and arises because leaving one participant out at a time changes both the slope and the intercept of the regression line. However, this variation is minuscule (the standard deviation is much smaller than the mean), indicating that our sample was sufficiently large. When pupil size was not corrected for luminosity, the average values of the coefficients were = 0.36723 0.06341, = 3.8296 0.0120, and the fit was inferior ( = 0.023 0.007).
3.3 Testing our model with ground truth obtained from independent judges rather than being self-reported
@@
We investigated how important it is to use self-reported arousal as ground truth. To do this, we used the approach other researchers have used: asking 10 independent judges (new participants) to give an arousal value to each video (e.g., Raiturkar et al. Raiturkar et al. (2016)). For each clip, we took the average value. We then used that arousal value in our method, instead of the self-reported one, to observe how much the average arousal of the independent judges could be predicted by pupil size corrected for the luminosity of our 47 participants. The comparison between the results obtained with our model using as arousal values those self-reported by the 47 participants, i.e., what we described so far, and the results obtained with our model using as arousal values those given by the 10 independent judges, is shown in the last column of the Table 8. The model trained on the arousal reported by the 10 independent judges provides clearly worse results than the one trained on the arousal reported by the 47 participants. For example, there is a worsening in of 188%. This is because when training a model, one must use a ground truth recorded from the same subjects from which the predictors are recorded, in this case, respectively, the self-reported arousal and pupil size of the 47 participants in the study. The reason is that the biosignals recorded from a given participant convey information about that participant’s emotional state.
@@
| Metric | Our Model | Hyperbolic Model Nakayama et al. (2021) | Linear Model Raiturkar et al. (2016); Asano et al. (2021a) | Our Model (without self-reported arousal) Raiturkar et al. (2016) |